OpenStudy (anonymous):

Help please :D !!!!!!!! Fayed tipped the pizza delivery driver three dollars for pizzas that cost eighteen dollars. Approximately what percent did he tip the driver? Round your answer to the nearest percent.

OpenStudy (anonymous):

A. two percent B. nine percent C. eleven percent D. sixteen percent

OpenStudy (ranga):

Percentage tipped = Amount tipped / cost * 100 = 3 / 18 * 100 = ?

OpenStudy (anonymous):

D

OpenStudy (anonymous):

@Niall @HeatDude

OpenStudy (anonymous):

@KattGirl14

OpenStudy (anonymous):

D :)

OpenStudy (austinl):

CashBoy, ranga gave you how to get the answer. All you need to do is type it into a calculator bro.

OpenStudy (anonymous):

@Darrius @GodIsMySavior @msumner @Confucious @thomaster @Isaiah.Feynman

OpenStudy (anonymous):

U ppl sure?

OpenStudy (anonymous):

i told you the answer it is D.3/18= .16666666 etc.... but if you make it into a a percentage its 16%

OpenStudy (anonymous):

ok I trust u ppl